# µOpTime: Statically Reducing the Execution Time of Microbenchmark Suites Using Stability Metrics

### v1

@@ 1. Introduction
We apply our approach, as well as two baseline approaches (i.e., minimum and random repetition levels), to three Go projects and eleven Java projects with $4,304$ microbenchmarks in total. Some Java projects originate from two different research data sets, which is why we introduce the terms data set 1 for the Go projects, and data set 2 and data set 3 for the two Java data sets. Crucially, the benchmarking results in data set 2 do not use warmup iterations in their result data, so that we can study the influence later (cf. RQ3). We observe that, compared to the baseline approaches, µOpTime provides better result accuracy, and handles the trade-off between execution time and accuracy better. Using µOpTime, we can reduce the execution time by up to $95.83\%$ for Go (median of $76.8\%$), by up to $35.96\%$ for Java projects in data set 2 (median of $13.08\%$), and by up to $71.74\%$ for Java projects in data set 3 (median of $26.78\%$), where at least 80% of the microbenchmarks being within 3% of their original result with the full configuration. Some configurations of µOpTime lead to higher time reductions, but especially for Java, this leads to lower accuracy (details in Section 4.2). µOpTime shortened the duration from $57.5$ to $2.32$ minutes for zap, the Go project with the largest absolute reduction, and from $178.19$ to $114.11$ hours for rxjava, the Java project from data set 2 with the largest absolute reduction. The Java project from data set 3 with the largest absolute time reduction was xodus with a reduction of $35.05$ minutes. The median execution time reductions were $33.6$ minutes for Go, $2.45$ hours for Java in data set 2, and $17.45$ minutes for Java in data set 3. For some stability metrics, such as relative median absolute deviation, the result quality of Java projects is negatively impacted, though, as more than $50\%$ of microbenchmarks have results deviating more than $3\%$ from the baseline result.
@@ 2. Background
RMIT (Randomized Multiple Interleaved Trials) is a randomization technique for eliminating and counteracting random fluctuation and measurement bias in microbenchmark results using multiple repetition levels (Abedi et al., 2015; Abedi and Brecht, 2017). Microbenchmarks in particular often run faster on repeated runs, because the underlying operating system and hardware use caching and speculative execution. To get statistically sound, comparable results, microbenchmarks within a suite have to be shuffled to minimize influence from speculative execution and caching, since following iterations of a particular microbenchmark are not executed directly after one another. In our setup, a suite is executed on a number of virtual cloud instances ($v$), the full suite is executed on each instance $s$ times, each benchmark within a suite is repeatedly executed for a number of iterations ($i$) following RMIT sequencing, and each iteration executes the evaluated function as often as possible for a fixed duration ($d$). In total, we define a tuple $(v,s,i,d)$ as the execution configuration for a microbenchmark. Each entry in this tuple also represents the different repetition levels. It is possible to implement different repetition levels to counteract other sources of measurement fluctuations. An example configuration can be seen in Figure 1. Using different repetition levels for microbenchmarks provides clear advantages for eliminating noise in measurements, but comes at the cost of increasing execution time and cost.
@@ 3.1. Optimization
Our optimization algorithm works with data that use a universal configuration of repetition levels, such as the levels in an RMIT setup, for all microbenchmarks. These levels are represented by a tuple $(n_{1},\dots,n_{l})\in\mathbb{N}^{+}$ with $l\in\mathbb{N}^{+}$ being the number of levels chosen for the experiment, and each $n_{i}$ being the number of repetitions inside a level. As an example, we might choose two levels for an experiment, where one level represents the iterations ($i$) of a benchmark, and the other level represents their duration ($d$). We then have to choose the maximum number of repetitions for each level in the experiment, such as $i=3$ iterations and $d=5$ seconds duration, making the tuple represent the levels $(3,5)$. [^1] To simplify the use of durations of RMIT setups here, we only consider full seconds of duration going forward. As a general rule, we order the tuple with the outermost level first. In our example, the outermost level is iterations (left), since each iteration runs for the duration (right).
@@ 22            $MIN\leftarrow MIN\cup S$
Algorithm 1 depicts our optimization algorithm. As input, it requires (i) the set of microbenchmarks within a suite $B$, (ii) the full execution configuration tuple $e_{max}$, (iii) the measurement values from the full suite execution $M$, (iv) a stability metric $sm$, and (v) a stability threshold value $ts$ for classifying microbenchmark configurations as (in)stable. (i)–(iii) are determined by the microbenchmarks, which a software project has implemented, and their settings for repetition levels. (iv) and (v) are settings for µOpTime itself, and adjust the behavior.
@@
The time complexity of Algorithm 1 is $O(|B|\times|E|)$, where $|B|$ is the amount of microbenchmarks, and $|E|$ is the amount of all execution configurations smaller than $e_{max}$. Since results of microbenchmarks do not depend on each other, the outer loop can be trivially parallelized. For the Go projects we investigate later in our evaluation, $|B|$ is between 20 and 40 (though projects exist with $|B|>1000$), while $|E|=3\times 5\times 5$, as we use an RMIT setup, where 3 repetition levels are optimized (see Section 4.1).
@@ 3.2. Integrating µOpTime in CI/CD pipelines
Figure 2 shows the integration of µOpTime into CI/CD pipelines. At first, because we do not know which microbenchmarks are more stable or unstable, they all have to be evaluated with the full configuration, which implies a long execution time. µOpTime then uses these results, analyzes which microbenchmarks are stable, and identifies for each microbenchmark an individual execution configuration which ensures reliable results while avoiding unnecessary executions. The individual and minimal execution configurations can then be used in regular CI/CD pipelines to quickly verify performance characteristics. The minimal configurations must be re-determined regularly, depending on how often and how severely the software is modified, for those microbenchmarks affected by software modifications.
@@ 3.3. Implementation
µOpTime is implemented in an open-source Python prototype, available in our replication package. [^2] Replication package: https://github.com/njapke/uoptime-replication-package It reads microbenchmarking results from a CSV file, identifies minimal execution configurations, and writes the results back into new CSV files. All five stability metrics from Section 2 (CV, RMAD, RCIW1, RCIW2, and RCIW3) are already implemented, adding additional ones is simple. The results contain the benchmarking score (mean or median performance depending on analyzed setup), CI, and a lower bound on time saved (theoretical calculation, not including overhead) for both the full and reduced execution configurations.
@@ 4. Experimental Evaluation
To evaluate our approach, we use two existing public data sets and collect a third one, studying 14 open source projects written in Go (3 projects) or Java (11 projects) (details in Section 4.1). We then explore our research questions RQ1 to RQ4 in Sections 4.2, 4.3, 4.4 and 4.5.
@@ 4.1. Data Sets
We use three different data sets to evaluate our approach. The data sets we use contain microbenchmarking results in different programming languages (Go, Java), since we aim to provide generalizable results. Furthermore, the two data sets containing Java microbenchmarking results are public data sets, and have been used in research before (Laaber et al., 2020b, 2021c, a, 2021b). Information on all data sets is summarized in Table 1.
@@
The first research question concerns the time savings of our approach. We apply µOpTime to all study subjects from all three data sets to get a minimal configuration for each microbenchmark. Then, we calculate the lowest possible time the minimal and full execution configurations take (i.e., the execution time without any overhead), and compare how much time the minimal configuration saves. For comparison, we include two simple baseline approaches, as other state-of-the-art approaches are orthogonal to µOpTime (Section 1). The first baseline approach (minimum baseline) always picks the smallest possible execution configuration, thereby maximizing time savings, while the second baseline approach (random baseline) picks a random execution configuration between the full and smallest execution configuration.
@@
To assess the effects on accuracy, we calculate a change rate $c$ for the performance results of each microbenchmark based on the results of the minimal and the full configuration ($r_{min}$ and $r_{full}$, respectively). The change rate is defined in Equation 1, and quantifies a relative change in measured results when using the minimal instead of the full configuration.
@@ Results
Figure 3 shows the results for all individual study subjects. The bars show a comparison between the execution time of the full and minimal configuration (with the time decreasing from the full bar to the smaller portion). The different marks show the fraction of microbenchmarks with a change rate in performance results of 1%, 3%, and 5%. We choose these different thresholds to better understand, where substantial differences in the results may be able to obfuscate potential performance regressions. This is informed by prior research, similar to how we set $ts$ (see Section 3.1), according to which performance measurements often vary by 3% (Georges et al., 2007), with performance regressions of 3% to 20% being relevant to practitioners (Huang et al., 2014). Thus, if a high number of microbenchmarks has change rates lower than our thresholds, the majority of relevant performance changes can still be detected.
The minimum baseline (Figures 3(a) and 3(b)) shows the highest possible time savings (95% to 99%) but has the lowest accuracy. It still fares well on Data set 1 (Go), with all projects having at least 62% of their microbenchmarks within 1% of the original result when comparing the mean, with that number increasing to 95% when comparing the median. For data set 2 (Java), the accuracy of the minimum baseline plummets, however, with all projects having at most 20% of their microbenchmarks with a 5% change or lower when comparing the mean. Using the median gives slightly better results, where at most 40% of each project’s microbenchmarks are within a 5% change of the original result. The result quality for data set 3 (Java) is between data set 1 and 2 for the minimum baseline, with all projects having at least 50% of their microbenchmarks with a 3% change or lower.
The random baseline (Figures 3(c) and 3(d)) shows time savings between 50% to 60% for data set 1 (Go), 70% to 75% for data set 2 (Java), and around 66% for data set 3 (Java). The accuracy loss is miniscule for data set 1, with the worst performing project still having 74.3% of microbenchmarks with a less than 1% change when comparing the mean, which jumps to 89.1% when comparing the median. For the java projects, the accuracy loss slightly higher. For data set 2 (java), at least 41% of each project’s microbenchmarks have a 3% change or lower when comparing the mean, with that number increasing to 64.1% when comparing the median. The results on data set 3 (Java) are better than on data set 2, with at least 68.9% of each project’s microbenchmarks showing a 3% change or lower when comparing the mean, with that number increasing to 74.3% when comparing the median.
@@ 4.3. RQ2: How does the choice of the stability metric affect the execution time reduction?
For each statistically different pair of stability metrics, we also calculate Cliff’s $\delta$ (Cliff, 1996) to quantify how strongly they differ. If the observations of two groups are the same, then $\delta=0$. Should group 1 be larger than group 2, then $\delta>0$, otherwise $\delta<0$. Cliff’s $\delta$ is bounded by $|\delta|\leq 1$. We also use the following commonly used categories (Romano et al., 2006): negligible ($|\delta|<0.147$), small ($0.330.147\leq|\delta|<$ 0.33), medium ($0.4740.33\leq|\delta|<$ 0.474), large ($|\delta|\geq 0.474$).
@@ Results
We conclude that the choice of stability metric largely does not matter for the software projects written in Go, as it has no statistically significant impact for all the software projects in data set 1. For the Java projects, RMAD produces consistently statistically different results, which, together with the overall lower accuracy (see Section 4.2), leads us to conclude that RMAD leads to worse results. In the case of rxjava, even more differences are statistically significant, but only RMAD has differences larger than “small” to the other stability metrics. To further illustrate this point using the data from RQ1, using RMAD for rxjava leads to 64.1% of microbenchmarks to deviate more than 10% from their original result. Even for data set 3, where the Kruskal-Wallis H test did not indicate statistically significant differences between stability metrics, using RMAD still produces higher change rates in most projects.
@@
Figure 4 summarizes the performance of µOpTime using RCIW3 with such a warmup phase. Counter-intuitively, time savings are considerably higher than without using a warmup phase, ranging from 12.14% to 63.58% of overall execution time, with median time savings of 42.77%. Accuracy is high, with all projects having more than 80% of microbenchmarks with a change rate smaller than 3% (median of 92.96%). This behavior is explained by the warmup behavior of the JVM. The JVM dynamically optimizes running code, leading to worse performance in the beginning, which then steadily becomes better, until it reaches a steady state (Traini et al., 2023). When we did not consider warmup, even iterations, which had not reached the steady state yet, were part of the measurement phase, increasing the overall instability of the microbenchmark (i.e., increasing the chosen stability metric). As these iterations are now part of the discarded warmup iterations, the instability is much lower, since only steady state measurements are used for calculating the stability. This leads to µOpTime finding smaller execution configurations, even though the number of warmup iterations per fork remain fixed. Therefore, we conclude that µOpTime performs better when considering warmup phases.
@@
We investigate to what extent the minimal configuration from µOpTime, minimum baseline, and random baseline (introduced in Section 4.2) detect the same performance changes as the full configuration. For all projects in data set 1 and 3, we compare all available versions in sequence, i.e., we compare $v_{1}$ to $v_{2}$, $v_{2}$ to $v_{3}$, and so on. The available versions are listed in Table 1 in Section 4.1. For some projects, we were unable to compare microbenchmarks beyond a certain version, because the developers replaced all microbenchmarks, which gives us a smaller number of usable version pairs for some projects. Our ground truth are the relevant performance changes detected by the full configuration. For each microbenchmark, we note whether the minimal configuration reports a true/false positive/negative, and calculate the false positive rate (FPR) and false negative rate (FNR). Since we always compare µOpTime and the two baselines to the full execution using the same underlying software versions and data, any code changes affect both the minimal and full execution.
@@
Table 2 shows the results for data set 1, and Table 3 shows the results for data set 3. The minimized configurations from µOpTime perform exceptionally in detecting the relevant performance changes in most cases.
For the Go projects, the minimized configurations using µOpTime identify 55 out of 57 performance changes correctly, and wrongly detect or miss performance changes for just 3 out of 12 version pairs. The FPR never exceeds 5%, while the FNR is at most 8.33%, with just 3 false positives and 2 false negatives in 348 microbenchmark comparisons. The baselines perform similarly well, with the minimum baseline having the same number of false positives and negatives, just at different microbenchmarks. The random baseline performs the best, with only 1 false positive and 1 false negative. Considering the time savings reported in Section 4.2, this is not too surprising, as the stability metric we use here, RMAD, has almost the same time savings as the minimum baseline, while the random baseline takes more time. The difference in quality inside a CI/CD pipeline is almost negligible as well, therefore we conclude that µOpTime works for Go projects, although our baselines perform similarly well.

### v2

@@ Abstract.
ACM Reference format:
Nils Japke, Martin Grambow, Christoph Laaber, and David Bermbach. 2025. µOpTime: Statically Reducing the Execution Time of Microbenchmark Suites Using Stability Metrics. ACM Trans. Softw. Eng. Methodol. 34, 8, Article 231 (October 2025), 26 pages.
https://doi.org/10.1145/3715322
@@ 1. Introduction
We apply our approach, as well as two baseline approaches (i.e., minimum and random repetition levels), to three Go projects and eleven Java projects with $4,304$ microbenchmarks in total. Some Java projects originate from two different research data sets, which is why we introduce the terms data set 1 for the Go projects, and data set 2 and data set 3 for the two Java data sets. Crucially, the benchmarking results in data set 2 do not use warmup iterations in their result data, so that we can study the influence later (cf. RQ3). We observe that, compared to the baseline approaches, µOpTime provides better result accuracy, and handles the trade-off between execution time and accuracy better. Using µOpTime, we can reduce the execution time by up to $95.83\%$ for Go (median of $76.8\%$), by up to $35.96\%$ for Java projects in data set 2 (median of $13.08\%$), and by up to $71.74\%$ for Java projects in data set 3 (median of $26.78\%$), where at least 80% of the microbenchmarks being within 3% of their original result with the full configuration. Some configurations of µOpTime lead to higher time reductions, but especially for Java, this leads to lower accuracy (details in Section˜4.2). µOpTime shortened the duration from $57.5$ to $2.32$ minutes for zap, the Go project with the largest absolute reduction, and from $178.19$ to $114.11$ hours for rxjava, the Java project from data set 2 with the largest absolute reduction. The Java project from data set 3 with the largest absolute time reduction was xodus with a reduction of $35.05$ minutes. The median execution time reductions were $33.6$ minutes for Go, $2.45$ hours for Java in data set 2, and $17.45$ minutes for Java in data set 3. For some stability metrics, such as relative median absolute deviation, the result quality of Java projects is negatively impacted, though, as more than $50\%$ of microbenchmarks have results deviating more than $3\%$ from the baseline result.
@@ 2. Background
RMIT (Randomized Multiple Interleaved Trials) is a randomization technique for eliminating and counteracting random fluctuation and measurement bias in microbenchmark results using multiple repetition levels (Abedi et al., 2015; Abedi and Brecht, 2017). Microbenchmarks in particular often run faster on repeated runs, because the underlying operating system and hardware use caching and speculative execution. To get statistically sound, comparable results, microbenchmarks within a suite have to be shuffled to minimize influence from speculative execution and caching, since following iterations of a particular microbenchmark are not executed directly after one another. In our setup, a suite is executed on a number of virtual cloud instances ($v$), the full suite is executed on each instance $s$ times, each benchmark within a suite is repeatedly executed for a number of iterations ($i$) following RMIT sequencing, and each iteration executes the evaluated function as often as possible for a fixed duration ($d$). In total, we define a tuple $(v,s,i,d)$ as the execution configuration for a microbenchmark. Each entry in this tuple also represents the different repetition levels. It is possible to implement different repetition levels to counteract other sources of measurement fluctuations. An example configuration can be seen in Figure˜1. Using different repetition levels for microbenchmarks provides clear advantages for eliminating noise in measurements, but comes at the cost of increasing execution time and cost.
@@ 3.1. Optimization
Our optimization algorithm works with data that use a universal configuration of repetition levels, such as the levels in an RMIT setup, for all microbenchmarks. These levels are represented by a tuple $(n_{1},\dots,n_{l})\in\mathbb{N}^{+}$ with $l\in\mathbb{N}^{+}$ being the number of levels chosen for the experiment, and each $n_{i}$ being the number of repetitions inside a level. As an example, we might choose two levels for an experiment, where one level represents the iterations ($i$) of a benchmark, and the other level represents their duration ($d$). We then have to choose the maximum number of repetitions for each level in the experiment, such as $i=3$ iterations and $d=5$ seconds duration, making the tuple represent the levels (3,5)(3,5). [^1] To simplify the use of durations of RMIT setups here, we only consider full seconds of duration going forward. As a general rule, we order the tuple with the outermost level first. In our example, the outermost level is iterations (left), since each iteration runs for the duration (right).
@@ 22    $MIN\leftarrow MIN\cup S$
Algorithm˜1 depicts our optimization algorithm. As input, it requires (i) the set of microbenchmarks within a suite $B$, (ii) the full execution configuration tuple $e_{max}$, (iii) the measurement values from the full suite execution $M$, (iv) a stability metric $sm$, and (v) a stability threshold value $ts$ for classifying microbenchmark configurations as (in)stable. (i)–(iii) are determined by the microbenchmarks, which a software project has implemented, and their settings for repetition levels. (iv) and (v) are settings for µOpTime itself, and adjust the behavior.
@@
The time complexity of Algorithm˜1 is $O(|B|\times|E|)$, where $|B|$ is the amount of microbenchmarks, and $|E|$ is the amount of all execution configurations smaller than $e_{max}$. Since results of microbenchmarks do not depend on each other, the outer loop can be trivially parallelized. For the Go projects we investigate later in our evaluation, $|B|$ is between 20 and 40 (though projects exist with $|B|>1000$), while $|E|=3\times 5\times 5$, as we use an RMIT setup, where 3 repetition levels are optimized (see Section˜4.1).
@@ 3.2. Integrating µOpTime in CI/CD pipelines
Figure˜2 shows the integration of µOpTime into CI/CD pipelines. At first, because we do not know which microbenchmarks are more stable or unstable, they all have to be evaluated with the full configuration, which implies a long execution time. µOpTime then uses these results, analyzes which microbenchmarks are stable, and identifies for each microbenchmark an individual execution configuration which ensures reliable results while avoiding unnecessary executions. The individual and minimal execution configurations can then be used in regular CI/CD pipelines to quickly verify performance characteristics. The minimal configurations must be re-determined regularly, depending on how often and how severely the software is modified, for those microbenchmarks affected by software modifications.
@@ 3.3. Implementation
µOpTime is implemented in an open-source Python prototype, available in our replication package (Japke et al., 2025). It reads microbenchmarking results from a CSV file, identifies minimal execution configurations, and writes the results back into new CSV files. All five stability metrics from Section˜2 (CV, RMAD, RCIW1, RCIW2, and RCIW3) are already implemented, adding additional ones is simple. The results contain the benchmarking score (mean or median performance depending on analyzed setup), CI, and a lower bound on time saved (theoretical calculation, not including overhead) for both the full and reduced execution configurations.
@@ 4. Experimental Evaluation
To evaluate our approach, we use two existing public data sets and collect a third one, studying 14 open source projects written in Go (3 projects) or Java (11 projects) (details in Section˜4.1). We then explore our research questions RQ1 to RQ4 in Sections˜4.2, 4.3, 4.4 and 4.5.
@@ 4.1. Data Sets
We use three different data sets to evaluate our approach. The data sets we use contain microbenchmarking results in different programming languages (Go, Java), since we aim to provide generalizable results. Furthermore, the two data sets containing Java microbenchmarking results are public data sets, and have been used in research before (Laaber et al., 2020b, 2021c, a, 2021b). Information on all data sets is summarized in Table˜1.
@@
  • a
    All projects are hosted on GitHub, except where noted.
  • b
    Only the newest version is listed. For commit hashes of all versions of data set 1, see our replication package (Japke et al., 2025), for data set 2 and 3, see the respective replication package (Laaber et al., 2020b, 2021c). Other versions than the newest one are only used for analyzing performance changes in RQ4 (Section˜4.5).
  • c
    Whenever the amount of microbenchmarks changes between versions we used, the range is indicated.
  • d
    Module directory in repository: https://github.com/openjdk/jmh
@@
The first research question concerns the time savings of our approach. We apply µOpTime to all study subjects from all three data sets to get a minimal configuration for each microbenchmark. Then, we calculate the lowest possible time the minimal and full execution configurations take (i.e., the execution time without any overhead), and compare how much time the minimal configuration saves. For comparison, we include two simple baseline approaches, as other state-of-the-art approaches are orthogonal to µOpTime (Section˜1). The first baseline approach (minimum baseline) always picks the smallest possible execution configuration, thereby maximizing time savings, while the second baseline approach (random baseline) picks a random execution configuration between the full and smallest execution configuration.
@@
To assess the effects on accuracy, we calculate a change rate $c$ for the performance results of each microbenchmark based on the results of the minimal and the full configuration ($r_{min}$ and $r_{full}$, respectively). The change rate is defined in Equation˜1, and quantifies a relative change in measured results when using the minimal instead of the full configuration.
@@ Results
Figure˜3 shows the results for all individual study subjects. The bars show a comparison between the execution time of the full and minimal configuration (with the time decreasing from the full bar to the smaller portion). The different marks show the fraction of microbenchmarks with a change rate in performance results of 1%, 3%, and 5%. We choose these different thresholds to better understand, where substantial differences in the results may be able to obfuscate potential performance regressions. This is informed by prior research, similar to how we set $ts$ (see Section˜3.1), according to which performance measurements often vary by 3% (Georges et al., 2007), with performance regressions of 3% to 20% being relevant to practitioners (Huang et al., 2014). Thus, if a high number of microbenchmarks has change rates lower than our thresholds, the majority of relevant performance changes can still be detected.
The minimum baseline (Figures˜3(a) and 3(b)) shows the highest possible time savings (95% to 99%) but has the lowest accuracy. It still fares well on Data set 1 (Go), with all projects having at least 62% of their microbenchmarks within 1% of the original result when comparing the mean, with that number increasing to 95% when comparing the median. For data set 2 (Java), the accuracy of the minimum baseline plummets, however, with all projects having at most 20% of their microbenchmarks with a 5% change or lower when comparing the mean. Using the median gives slightly better results, where at most 40% of each project’s microbenchmarks are within a 5% change of the original result. The result quality for data set 3 (Java) is between data set 1 and 2 for the minimum baseline, with all projects having at least 50% of their microbenchmarks with a 3% change or lower.
The random baseline (Figures˜3(c) and 3(d)) shows time savings between 50% to 60% for data set 1 (Go), 70% to 75% for data set 2 (Java), and around 66% for data set 3 (Java). The accuracy loss is miniscule for data set 1, with the worst performing project still having 74.3% of microbenchmarks with a less than 1% change when comparing the mean, which jumps to 89.1% when comparing the median. For the java projects, the accuracy loss slightly higher. For data set 2 (java), at least 41% of each project’s microbenchmarks have a 3% change or lower when comparing the mean, with that number increasing to 64.1% when comparing the median. The results on data set 3 (Java) are better than on data set 2, with at least 68.9% of each project’s microbenchmarks showing a 3% change or lower when comparing the mean, with that number increasing to 74.3% when comparing the median.
@@ 4.3. RQ2: How does the choice of the stability metric affect the execution time reduction?
For each statistically different pair of stability metrics, we also calculate Cliff’s $\delta$ (Cliff, 1996) to quantify how strongly they differ. If the observations of two groups are the same, then $\delta=0$. Should group 1 be larger than group 2, then $\delta>0$, otherwise $\delta<0$. Cliff’s $\delta$ is bounded by $|\delta|\leq 1$. We also use the following commonly used categories (Romano et al., 2006): negligible ($|\delta|<0.147$), small ($0.147\leq|\delta|<0.33$), medium ($0.33\leq|\delta|<0.474$), large ($|\delta|\geq 0.474$).
@@ Results
We conclude that the choice of stability metric largely does not matter for the software projects written in Go, as it has no statistically significant impact for all the software projects in data set 1. For the Java projects, RMAD produces consistently statistically different results, which, together with the overall lower accuracy (see Section˜4.2), leads us to conclude that RMAD leads to worse results. In the case of rxjava, even more differences are statistically significant, but only RMAD has differences larger than “small” to the other stability metrics. To further illustrate this point using the data from RQ1, using RMAD for rxjava leads to 64.1% of microbenchmarks to deviate more than 10% from their original result. Even for data set 3, where the Kruskal-Wallis H test did not indicate statistically significant differences between stability metrics, using RMAD still produces higher change rates in most projects.
@@
Figure˜4 summarizes the performance of µOpTime using RCIW3 with such a warmup phase. Counter-intuitively, time savings are considerably higher than without using a warmup phase, ranging from 12.14% to 63.58% of overall execution time, with median time savings of 42.77%. Accuracy is high, with all projects having more than 80% of microbenchmarks with a change rate smaller than 3% (median of 92.96%). This behavior is explained by the warmup behavior of the JVM. The JVM dynamically optimizes running code, leading to worse performance in the beginning, which then steadily becomes better, until it reaches a steady state (Traini et al., 2023). When we did not consider warmup, even iterations, which had not reached the steady state yet, were part of the measurement phase, increasing the overall instability of the microbenchmark (i.e., increasing the chosen stability metric). As these iterations are now part of the discarded warmup iterations, the instability is much lower, since only steady state measurements are used for calculating the stability. This leads to µOpTime finding smaller execution configurations, even though the number of warmup iterations per fork remain fixed. Therefore, we conclude that µOpTime performs better when considering warmup phases.
@@
We investigate to what extent the minimal configuration from µOpTime, minimum baseline, and random baseline (introduced in Section˜4.2) detect the same performance changes as the full configuration. For all projects in data set 1 and 3, we compare all available versions in sequence, i.e., we compare $v_{1}$ to $v_{2}$, $v_{2}$ to $v_{3}$, and so on. The available versions are listed in Table˜1 in Section˜4.1. For some projects, we were unable to compare microbenchmarks beyond a certain version, because the developers replaced all microbenchmarks, which gives us a smaller number of usable version pairs for some projects. Our ground truth are the relevant performance changes detected by the full configuration. For each microbenchmark, we note whether the minimal configuration reports a true/false positive/negative, and calculate the false positive rate (FPR) and false negative rate (FNR). Since we always compare µOpTime and the two baselines to the full execution using the same underlying software versions and data, any code changes affect both the minimal and full execution.
@@
Table˜2 shows the results for data set 1, and Table˜3 shows the results for data set 3. The minimized configurations from µOpTime perform exceptionally in detecting the relevant performance changes in most cases.
For the Go projects, the minimized configurations using µOpTime identify 55 out of 57 performance changes correctly, and wrongly detect or miss performance changes for just 3 out of 12 version pairs. The FPR never exceeds 5%, while the FNR is at most 8.33%, with just 3 false positives and 2 false negatives in 348 microbenchmark comparisons. The baselines perform similarly well, with the minimum baseline having the same number of false positives and negatives, just at different microbenchmarks. The random baseline performs the best, with only 1 false positive and 1 false negative. Considering the time savings reported in Section˜4.2, this is not too surprising, as the stability metric we use here, RMAD, has almost the same time savings as the minimum baseline, while the random baseline takes more time. The difference in quality inside a CI/CD pipeline is almost negligible as well, therefore we conclude that µOpTime works for Go projects, although our baselines perform similarly well.